0 0
Add a comment Improve this question Transcribed image text
Answer #1

a)

Time Constant

T=RC =(4*10^6)(2*10^-6)

T=8 sec

In a RC Circuit voltage as a function of time is given by

V=Vo[1-e-t/T]

V=16[1-e-5/8]

V=7.44 Volts

Voltage drop across resistor is

VR=16-7.44=8.56 Volts

b)

Time Constant

T=Req*C =[(4+7)*106](2*10-6]

T=22 sec

For discharging ,Voltage drop across both the resistors is

V=Voe-t/T

V=16*e-(1.8/22)

V=14.74 Volts

Voltage across R1 is

V1=V*(R1/R1+R2) =14.74*(4/4+7)

V1=5.36 Volts